A toroidal solenoid has 3000 turns and a radius of 2cm with a current of 5A flowing through it. Determine the magnitude of the magnetic field.


1
Expert's answer
2021-05-13T14:26:05-0400

"B=\frac{\mu_0 nI}{2\pi r}=\frac{2\cdot 10^{-7}\cdot 3000\cdot 5}{0.02}=0.15~T."